What is the sum the interior angles of a 13-gon?
What is the sum the interior angles of a 20-gon?
What is the sum the interior angles of a 18-gon?
Answer #1
The sum of interior angle ( Sum-of-interior-angles-n-gon) of n-gon is: {{{Sum-of-interior-angles-n-gon = (n-2)180}}} If we have a {{{13-gon}}}, {{{n=13}}} Then {{{Sum-of-interior-angles-n-gon = (13-2)180}}} {{{Sum-of-interior-angles-n-gon = (11)180}}} {{{Sum-of-interior-angles-n-gon = (11)180}}} {{{Sum-of-interior-angles-n-gon = 1980}}} If we have {{{20-gon}}}, {{{n=20}}} {{{Sum-of-interior-angles-n-gon = (20-2)180}}} {{{Sum-of-interior-angles-n-gon = (18)180}}} {{{Sum-of-interior-angles-n-gon = 3240}}} If we have {{{18-gon}}}, {{{n=18}}} {{{Sum-of-interior-angles-n-gon = (18-2)180}}} {{{Sum-of-interior-angles-n-gon = (16)180}}} {{{Sum-of-interior-angles-n-gon = 2880}}}
answered by: Acorn